OpenStudy (anonymous):

Solve the system: 3x-y=-1 2x-y+z=-6 x+4y-z=9

OpenStudy (candyme):

\[x+4y-z=9\]\[3x-y=-12\]\[x-y+z=-6\] Is this system correct?

OpenStudy (anonymous):

no. 3x-y=-1 2x-y+z=-6 x+4y-z=9

OpenStudy (saifoo.khan):

Step1: Take any equation. Let's say 2. Solve it for any variable. Let's say z. It will look like: z = -6 + y - 2x Now take eq 3. Solve it for z. z = x + 4y + 9

OpenStudy (saifoo.khan):

Final thing. eq 2 = eq 3 -6 + y - 2x = x + 4y + 9 Simplify.

OpenStudy (anonymous):

with this you add all the like factors right?

OpenStudy (saifoo.khan):

You will get something like equation 1. Then solving that equation and equation 1 simultaneously will give you value of x and y

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

and i have 2 more problems like this

OpenStudy (saifoo.khan):

-6 + y - 2x = x + 4y + 9 Simplify. -15 = 6y + 3y Do you get till here?

OpenStudy (anonymous):

how did you get 6y when it is 4y and a y wouldnt it be 5y and why is there two ys shouldnt there be an x so no i dont get it

OpenStudy (saifoo.khan):

Sorry. Did a typo. -6 + y - 2x = x + 4y + 9 Simplify. -15 = 3x + 3y Do you get till here?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

Now take this equation AND equation 1 -15 = 3x + 3y AND 3x-y=-1 Solve both of these.

OpenStudy (anonymous):

how? i have questions like that i need help with to

OpenStudy (saifoo.khan):

You can solve this by elimination or substitution.

OpenStudy (anonymous):

i know that i get confused doing it.. I did a couple but then some i got confused on can give me the steps please on how to do it so that i can do it on the others i need help with and use that example

OpenStudy (anonymous):

to look back to if that makes since

OpenStudy (saifoo.khan):

Just out of curiosity, from how many decades you haven't been doing your homework? XD

OpenStudy (saifoo.khan):

Is your computer still locked?

OpenStudy (anonymous):

yes.. and what was your last question?

OpenStudy (saifoo.khan):

You have soo much homework! D: I can send you links for examples but your computer won't allow.. How can you use Openstudy?

OpenStudy (anonymous):

my mom unlocked it for me but she is busy and i now i have tons of homework it is because i just had surgery and need to catch up

OpenStudy (anonymous):

you can try and use the links and i can see if it works

OpenStudy (anonymous):

nope doesnt work

OpenStudy (anonymous):

that one worked

OpenStudy (saifoo.khan):

http://www.youtube.com/watch?v=uzyd_mIJaoc

OpenStudy (anonymous):

ok let me see if i did this right.. 15=3x+3y 1=3x-y correct

OpenStudy (anonymous):

then i multiplied -15=3x+3y by -1 so that the 3x would become a negative so that made it 15=-3x-3y so the x's cancelled out which made it 15=-3y 1=-y then that was 16=-4y then i added 4 on each side to get y by it self and that became 20 y is that right so far

OpenStudy (anonymous):

so y=20 so i plugged that in 15=3x+3(20) 1=3x-20 15=3(x)=60 1=3x-20 16=6x+40 then i subtracted 16 on each side so it became 6x+24 and i dont know wheather to subtract the 6 and get x=18 or divide and get x=4